OpenStudy (anonymous):

find mean value? 2cos2x+sinx=0 a= -pi\2, b=pi\2

OpenStudy (anonymous):

plz reply karo plzzzzzzzzzz

myininaya (myininaya):

{f(b)-f(a)}/{b-a}={f(pi/2)-f(-pi/2)}/{pi}={[2cos(pi)+sin(pi/2)]-[2cos(-pi)+sin(-pi/2)]}/{pi} ={2(-1)+1-2(-1)-(-1)}/pi=(-2+1+2+1)/pi=2/pi but f'(x)=4sin(2x)+cosx f'(c)=4sin(2c)+cos(c) set f'(c)=2/pi and solve for c

OpenStudy (anonymous):

i got a value of c

myininaya (myininaya):

ok cool!

OpenStudy (anonymous):

i got ! 4sin2c+(1\90)=cosc now how can be find value of c

myininaya (myininaya):

4sin(2c)+cos(c)=2/pi

OpenStudy (anonymous):

plz reply fast myninaya............

myininaya (myininaya):

-0.04535615498 is what I got using a calculator so the problem says to use the mean value thm?

OpenStudy (anonymous):

how can you find the value of c . give me the method?

OpenStudy (dumbcow):

\[Average = \frac{1}{b-a}\int\limits_{a}^{b}2\cos 2x +\sin x dx\]

OpenStudy (anonymous):

dumcow i don't understand this method . give me another method?

OpenStudy (dumbcow):

are you looking for the mean value of the function from a to b or does it say to use mean value theorem to find point that equals avg rate of change

OpenStudy (anonymous):

but my teacher says that u can use mean value theorm

myininaya (myininaya):

http://www.eldamar.org.uk/maths/calculus/node17.html

myininaya (myininaya):

m(b-a)=int(f(x),a..b) m is the mean value

myininaya (myininaya):

cow has already said this though

myininaya (myininaya):

do you know how to integrate?

OpenStudy (anonymous):

no i don't know?

OpenStudy (anonymous):

Okay, By the mean value theorem. The mean value is [f(b)-f(a)]/[b-a]. \[f(b)=f(\pi/2)=2\cos \pi +\sin(\pi/2)=-2+1=-1\] \[f(a)=f(-{\pi \over 2})=2\cos (-\pi)+\sin (-\pi/2)=-2-1=-3\]

OpenStudy (anonymous):

Therefore, \[{f(b)-f(a) \over b-a}={-1-(-3) \over {\pi \over 2}-{-\pi \over 2}}={2 \over \pi}\]

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

its right but how we can find the value of c?

OpenStudy (anonymous):

This value, we just found (2/pi) is equal to f'(c). You can use this relation to find c.

OpenStudy (anonymous):

f`c= -4sin2c+cosc

OpenStudy (anonymous):

-4sin2c+cosc=pi\2 so how we can find C?

OpenStudy (anonymous):

Solve the equation, you may get more than one value for c. Take only the value that is in the given interval (-pi/2,pi/2).

OpenStudy (anonymous):

how can it solve the equa? i don't understand this equation to find the value for C

myininaya (myininaya):

i would graph it and approximate the solution

myininaya (myininaya):

you could also use newton's method

OpenStudy (anonymous):

You never solved quadratic equation?! Hmm I think myininaya got a point. It's difficult to solve it using identities. Probably graphing is a good method.

OpenStudy (anonymous):

Just gimme a minute.

OpenStudy (anonymous):

BRB

OpenStudy (anonymous):

BRB what?

myininaya (myininaya):

means be right back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no body can solve this question?

OpenStudy (anonymous):

oops myininaya has already solved the problem. Sorry I didn't see that.

myininaya (myininaya):

lol

OpenStudy (anonymous):

you are give my proper metjhod to find the valuc of C?

OpenStudy (anonymous):

as myininaya, we can estimate the value by graphing.

OpenStudy (anonymous):

c will be around 1.66

OpenStudy (anonymous):

you said that C1.66 . how you can find tell me?

OpenStudy (anonymous):

Wait. this value of c is out side our interval. The value of c, that's in the interval is around 0.045

myininaya (myininaya):

look for x intercepts of f'(c)=2/pi

OpenStudy (anonymous):

\[f'(c)={f(b)-f(a) \over b-a}\] this is the formula.

OpenStudy (anonymous):

ok i know thic formula i completed.. -4sin2c+cosx=2\pi after what can i do i don't understand?

myininaya (myininaya):

this is what I got when I used newton's method to find c

OpenStudy (anonymous):

f(x)=2cos2x+sinx andf`(x)=-4sin2x+cosx

myininaya (myininaya):

oops i forgot about the 2/pi you try i have to go

OpenStudy (anonymous):

ok thanks!